OpenStudy (anonymous):

-x^2 -3x+3 ; get into vertex form

OpenStudy (anonymous):

= -1(x^2+3x + 9/4 )+3- 2 *9/4 = -1(x^2+3x+9/4)-3/2 i dont know what to do after that

OpenStudy (anonymous):

f(x) = a(x-h)^2+k

OpenStudy (anonymous):

yeah but i need to be vertex form